# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v5

@@ Abstract
Real-world knowledge graphs (KGs) contain not only standard triple-based facts, but also more complex, heterogeneous types of facts, such as hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts that imply relationships between facts. These richer forms of representation have attracted significant attention due to their enhanced expressiveness and capacity to model complex semantics in real-world scenarios. However, most existing studies suffer from two main limitations: (1) they typically focus on modeling only specific types of facts, thus making it difficult to generalize to real-world scenarios with multiple fact types; and (2) they struggle to achieve generalizable hierarchical (inter-fact and intra-fact) modeling due to the complexity of these representations. To overcome these limitations, we propose UniHR, a Unified Hierarchical Representation learning framework, which consists of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing both semantic information within individual facts and enriching the structural information between facts. To go beyond the unified method itself, we further explore the potential of unified representation in complex real-world scenarios, including joint modeling of multi-task, compositional and hybrid facts. Extensive experiments on 9 datasets across 5 types of KGs demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations.
@@ Introduction
Real-world large-scale knowledge graphs (KGs) such as Wikidata (Vrandečić and Krötzsch 2014) and DBpedia (Lehmann et al. 2015) have been widely applied in many areas including question answering (Kaiser, Saha Roy, and Weikum 2021) and natural language processing (Annervaz, Chowdhury, and Dukkipati 2018; Liu et al. 2025a). To faithfully represent complex real-world knowledge, these KGs usually incorporate not only standard triple-based facts, but also more complex and heterogeneous types of facts such as hyper-relational, temporal and nested facts.
Despite the simplicity of triple-based representation (i.e., $(head,relation,tail)$), such forms struggle to capture the complexity of real-world facts, e.g., “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry”. Consequently, recent studies (Xiong et al. 2023b) have focused on semantically richer beyond-triple facts, including: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These fact types allow for expression of complex semantics and revelation of relationships between facts. Thus in recent years, Hyper-relational KGs (HKG) (Luo et al. 2023), Temporal KGs (TKG) (Zhang et al. 2025a), and Nested factual KGs (NKG) (Li et al. 2025) attract wide research interests.
As shown in Figure 1, we find that existing research mainly suffers from two major limitations: (1) They fail to reflect real-world scenarios with multiple heterogeneous fact types (Xiong et al. 2023b), instead artificially dividing and only modeling a single KG type; (2) Earlier triple-based studies (Chen et al. 2021) have demonstrated the effectiveness of hierarchical fact semantic modeling (inter-fact and intra-fact). But due to the complexity of beyond-triple representations, they struggle to achieve comprehensive hierarchical semantic modeling, even generalizing to other fact types. Specifically, for HKGs, StarE (Galkin et al. 2020) customizes GNN to enhance inter-fact interactions, while GRAN (Wang et al. 2021) designs attention variants with edge-bias to capture intra-fact heterogeneity. For NKGs, NestE (Xiong et al. 2024) et al. NKG methods connect bi-level facts and only score intra-fact semantics. For TKGs, either by explicitly incorporating temporal information into the score function like GeomE+ (Xu et al. 2023), or by unfolding entity neighborhood subgraphs along temporal chain and capturing inter-fact semantics to model temporal information like ECEformer (Fang et al. 2024). Although advanced methods like HAHE (Luo et al. 2023) begin to capture hierarchical semantics for HKGs, their heterogeneity in representation limits their scalability to other fact types. Therefore, establishing a unified hierarchical representation learning method for real-world KG with multiple fact types is worth investigating.
@@
This section discusses the limitations of existing methods in hierarchical modeling. See Appendix C for more discussion.
@@
Early HKG methods typically focus on modeling either local or global information. Galkin et al. (2020) customize StarE based on CompGCN (Vashishth et al. 2019) for H-Facts to capture global structure information among H-Facts, demonstrating the importance of structure information in HKGs. GRAN (Wang et al. 2021) with edge-aware bias in attention (Vaswani et al. 2017) layer, HyNT (Chung, Lee, and Whang 2023) with qualifier-aware encoder, and ShrinkE with relation-specific box (Xiong et al. 2023a) all focus on modeling intra-fact semantic information. Recent advanced methods aim to comprehensively capture both inter-fact and intra-fact information. For example, HAHE (Luo et al. 2023) employs dual-graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling, achieving significant performance improvements. Similarly, HyperSAT (Wang, Chen, and Zhang 2025) accomplishes this through a combination of graph sampling and a key-value joint attention mechanism. We consider hierarchical modeling of KGs to be a promising direction, existing approaches are customized for HKG form and difficult to generalize to other types of facts.
@@
Chung et al. (Chung and Whang 2023) are the first to introduce nested facts to model relationships between facts, and also propose BiVE which bridges semantics between atomic facts and fact nodes via a simple MLP and scores both atomic facts and nested facts using quaternion-based KGE scoring functions like QuatE (Zhang et al. 2019) or BiQUE (Guo and Kok 2021). Based on BiVE, NestE (Xiong et al. 2024) represents fact nodes using a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss. GRADATE (Li et al. 2025) enhances entity and fact representation learning by mining latent intra-fact semantics. However, due to the complexity of NKG representations, existing methods have so far primarily focused on capturing intra-fact semantic information.
@@
Recent studies on TKG representation learning have mainly focused on designing elegant time-aware modules to enhance representation capability. Advanced models like TGeomE+ (Xu et al. 2023) improve the modeling of local semantics in TKGs through 4th-order tensor factorization and linear temporal regularization. Similarly, HGE (Pan et al. 2024) and 5EL (Zhang et al. 2025a) enhance the expressiveness of temporal spaces by introducing geometric spaces. In contrast, ECEformer (Fang et al. 2024) captures only inter-fact semantics by unfolding entity neighborhood subgraphs along the temporal chain to model temporal information. However, TKG representation learning methods that can simultaneously capture both intra-fact temporal semantics and global structural information remain largely unexplored. UniHR achieves this by regarding timestamps as nodes and directly employing hierarchical message passing.
@@
Unlike existing triple-based systems (Ali et al. 2022) including RDF (triple) reification, RDF-star and labeled RDF representation techniques, we constrain “triple” to serve as the basic units of HiDR form, and split “nodes” and “relations” into three types respectively, making it more suitable for graph representation learning. Meanwhile, “triple” form makes HiDR could continuously benefit from the model developments of triple-based KGs, which is the most active area of research for link prediction over KGs.
@@
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec (Kazemi et al. 2019):
@@
NaLP (Guan et al. 2019) 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
StarE (Galkin et al. 2020) 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
GRAN (Wang et al. 2021) 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
tNaLP (Guan et al. 2023) 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
HyNT (Chung, Lee, and Whang 2023) 0.479 0.411 0.601 0.478 0.409 0.601 0.328 0.256 0.468 0.356 0.285 0.493
ShrinkE (Xiong et al. 2023a) 0.485 0.431 0.601 - - - 0.345 0.275 0.482 - - -
HAHE (Luo et al. 2023) 0.498 0.418 0.610 0.497 0.421 0.614 0.343 0.269 0.484 0.378 0.306 0.515
NYLON (Yu, Yang, and Yang 2024) 0.385 0.299 0.527 0.384 0.300 0.520 0.326 0.262 0.446 0.291 0.226 0.414
HyperSAT (Wang, Chen, and Zhang 2025) 0.493 0.427 0.610 0.496 0.430 0.613 0.345 0.270 0.489 0.380 0.306 0.520
@@
Table 1: Results on HKG datasets, are reproduced by us and others are taken from (Wang, Chen, and Zhang 2025).
@@
Since the query varies across different settings, we use the transformer (Vaswani et al. 2017) as the decoder with mask pattern. Specifically, we serialize the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
@@
For HKGs, we use WikiPeople (Guan et al. 2019) and WD50K (Wang et al. 2021). For NKGs, we select FBH, FBHE and DBHE (Chung and Whang 2023). For TKGs, we use wikidata12k (Dasgupta, Ray, and Talukdar 2018). To further evaluate the potential of the unified representation, we further introduce hyper-relational TKG datasets WIKI-hy and YAGO-hy (Ding et al. 2024). Detailed statics of 9 benchmarks above are given in Appendix E.
@@
We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings (Bordes et al. 2013) during the evaluation to eliminate existing facts in the dataset. It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set. Similarly, the difference in filtering settings of TKG occurs in timestamp. In this paper, we adopt the strict filtering setting of the former. To ensure fair comparison, for HKG we utilize the results of HyperSAT (Wang, Chen, and Zhang 2025) with the same settings as ours. For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
@@
For HKG, we compare with NaLP (Guan et al. 2019), StarE (Galkin et al. 2020), GRAN (Wang et al. 2021), tNaLP (Guan et al. 2023), HyNT (Chung, Lee, and Whang 2023), ShrinkE (Xiong et al. 2023a), HAHE (Luo et al. 2023), NYLON (Yu, Yang, and Yang 2024) and HyperSAT (Wang, Chen, and Zhang 2025). For NKG, QuatE (Zhang et al. 2019), BiQUE (Guo and Kok 2021), BiVE (Chung and Whang 2023), NestE (Xiong et al. 2024), HOKE (Pirrò 2025) and GRADATE (Li et al. 2025) are chosen as baselines. For TKG, we compare against following methods: ATiSE (Xu et al. 2019), TGeomE+ (Xu et al. 2023), HGE (Pan et al. 2024), HyperSAT (Wang, Chen, and Zhang 2025), DuaTHP (Chen et al. 2025), ECEformer (Fang et al. 2024) and 5EL (Zhang et al. 2025a).
@@
QuatE (Zhang et al. 2019) 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE (Guo and Kok 2021) 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
BiVE (Chung and Whang 2023) 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE (Xiong et al. 2024) 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
HOKE (Pirrò 2025) - - - - 3.06 0.719 0.777 2.82 0.674 0.764 2.10 0.674 0.777
GRADATE (Li et al. 2025) - - - - 18.15 0.780 0.871 26.81 0.603 0.757 4.72 0.654 0.916
@@
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from (Chung and Whang 2023) for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting. Hyperparameter details can be found in Appendix F.
@@
ATiSE (Xu et al. 2019) 0.252 0.148 0.288 0.462
TGeomE+ (Xu et al. 2023) 0.333 0.232 0.361 0.546
HGE (Pan et al. 2024) 0.290 0.176 0.323 0.514
ECEformer (Fang et al. 2024) 0.262 0.159 0.255 0.462
DuaTHP (Chen et al. 2025) 0.304 0.209 0.331 0.509
HyperSAT (Wang, Chen, and Zhang 2025) 0.313 0.240 0.348 0.456
5EL (Zhang et al. 2025a) 0.311 0.237 0.355 0.546
@@
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. Unlike previous methods that use customized decoders for triples, our unified approach does not. The effectiveness of UniHR with other decoders is shown in Appendix A.
@@
Figure 4: (a) improvements of joint training on hybrid tasks. (b) improvements of joint training on wikimix dataset with hybrid fact forms. Yellow region indicates improvements achieved by joint training. Detailed results are shown in Appendix D.
@@
To analyze the contribution of different modules across various KG types, we present ablation results in Table 4, with more studies provided in Appendix B. It can be observed that both intra-fact and inter-fact message passing contribute to performance improvement. In particular, intra-fact message passing proves to be more beneficial for NKGs. We attribute this to fact nodes in NKGs being inherently composed of other atomic nodes, making triple prediction rely heavily on comprehensive bi-level fact semantics. In contrast, HKGs and TKGs focus solely on atomic nodes, whose representations are not dependent on other nodes. Therefore, inter-fact message passing, by capturing the global context among facts, works more effective for HKGs and TKGs, leading to further performance improvements.
@@
Owing to its unified representation, UniHR can flexibly generalize to compositional knowledge graphs, such as hyper-relational temporal KGs (HTKGs) (Ding et al. 2024), which integrate the characteristics of both HKGs and TKGs. In HTKGs, each hyper-relational fact is associated with a timestamp that explicitly indicates its temporal validity. As shown in Table 5, UniHR offers a performance improvement in link prediction tasks on HTKGs, outperforming both TKG-specific and HKG-specific models. This result illustrates the strong ability of UniHR to jointly model auxiliary key-value pairs and temporal information. Furthermore, UniHR achieves competitive performance with the specialized model HypeTKG, despite not relying on complex module stacking.
@@
For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form. Therefore, we attempt joint training on two tasks using the NKG dataset, as shown in Figure 4(b). Consistent with previous studies (Li et al. 2025), we also observe that results of joint training are generally superior to those of separate training, further confirming that nested and atomic facts can mutually enhance and complement each other’s semantics.
@@
We believe unified representation is key to develop pre-trained models that integrate multiple KG types. To explore this potential, we jointly train on different KG types. Notably, real-world KGs like Wikidata (Vrandečić and Krötzsch 2014) naturally contain diverse fact types. Thus, we construct a hybrid dataset wikimix, by filtering two Wikidata subsets: HKG WikiPeople and TKG wikidata12k, which share 3,546 entities and 18 relations but have no overlapping facts. To prevent data leakage, we remove test entries whose main triples appear in the other subset’s train set (537 from wikidata12k and 384 from WikiPeople). Details are in Appendix E.

### v6

@@ Abstract
Real-world knowledge graphs (KGs) contain not only standard triple-based facts, but also more complex, heterogeneous types of facts, such as hyper-relational facts with auxiliary key-value pairs, temporal facts with additional timestamps, and nested facts that imply relationships between facts. These richer forms of representation have attracted significant attention due to their enhanced expressiveness and capacity to model complex semantics in real-world scenarios. However, most existing studies suffer from two main limitations: (1) they typically focus on modeling only specific types of facts, thus making it difficult to generalize to real-world scenarios with multiple fact types; and (2) they struggle to achieve generalizable hierarchical (inter-fact and intra-fact) modeling due to the complexity of these representations. To overcome these limitations, we propose UniHR, a Unified Hierarchical Representation learning framework, which consists of a learning-optimized Hierarchical Data Representation (HiDR) module and a unified Hierarchical Structure Learning (HiSL) module. The HiDR module unifies hyper-relational KGs, temporal KGs, and nested factual KGs into triple-based representations. Then HiSL incorporates intra-fact and inter-fact message passing, focusing on enhancing both semantic information within individual facts and enriching the structural information between facts. To go beyond the unified method itself, we further explore the potential of unified representation in complex real-world scenarios. Extensive experiments on 9 datasets across 5 types of KGs demonstrate the effectiveness of UniHR and highlight the strong potential of unified representations. Code and data are available at https://github.com/zjukg/UniHR.
@@ Introduction
Real-world large-scale knowledge graphs (KGs) such as Wikidata (wikidata) and DBpedia (DBpedia) have been widely applied in many areas including question answering (kaiser2021reinforcement) and natural language processing (annervaz2018learning; ontotune). To faithfully represent complex real-world knowledge, these KGs usually incorporate not only standard triple-based facts, but also more complex and heterogeneous types of facts such as hyper-relational, temporal and nested facts.
Despite the simplicity of triple-based representation (i.e., $(head,relation,tail)$), such forms struggle to capture the complexity of real-world facts, e.g., “Oppenheimer is educated at Harvard University for a bachelor degree in chemistry”. Consequently, recent studies (beyond_triple) have focused on semantically richer beyond-triple facts, including: hyper-relational fact ((Oppenheimer, educated at, Harvard University), degree: bachelor, major: chemistry), temporal fact (Oppenheimer, honored with, Fermi Prize, 1963), nested fact ((Oppenheimer, born in, New York), imply, (Oppenheimer, nationality, The United States)). These fact types allow for expression of complex semantics and revelation of relationships between facts. Thus in recent years, Hyper-relational KGs (HKG) (hahe), Temporal KGs (TKG) (5EL), and Nested factual KGs (NKG) (GRADATE) attract wide research interests.
As shown in Figure 1, we find that existing research mainly suffers from two major limitations: (1) They fail to reflect real-world scenarios with multiple heterogeneous fact types (beyond_triple), instead artificially dividing and only modeling a single KG type; (2) Earlier triple-based studies (hitter) have demonstrated the effectiveness of hierarchical fact semantic modeling (inter-fact and intra-fact). But due to the complexity of beyond-triple representations, they struggle to achieve comprehensive hierarchical semantic modeling, even generalizing to other fact types. Specifically, for HKGs, StarE (stare) customizes GNN to enhance inter-fact interactions, while GRAN (gran) designs attention variants with edge-bias to capture intra-fact heterogeneity. For NKGs, NestE (neste) et al. NKG methods connect bi-level facts and only score intra-fact semantics. For TKGs, either by explicitly incorporating temporal information into the score function like GeomE+ (TGeomE+), or by unfolding entity neighborhood subgraphs along temporal chain and capturing inter-fact semantics to model temporal information like ECEformer (eceformer). Although advanced methods like HAHE (hahe) begin to capture hierarchical semantics for HKGs, their heterogeneity in representation limits their scalability to other fact types. Therefore, establishing a unified hierarchical representation learning method for real-world KG with multiple fact types is worth investigating.
@@
This section discusses the limitations of existing methods in hierarchical modeling.
@@
Early HKG methods typically focus on modeling either local or global information. stare customize StarE based on CompGCN (CompGCN) for H-Facts to capture global structure information among H-Facts, demonstrating the importance of structure information in HKGs. GRAN (gran) with edge-aware bias in attention (transformer) layer, HyNT (hynt) with qualifier-aware encoder, and ShrinkE with relation-specific box (shrinke) all focus on modeling intra-fact semantic information. Recent advanced methods aim to comprehensively capture both inter-fact and intra-fact information. For example, HAHE (hahe) employs dual-graph attention and edge-aware bias in the transformer attention layer for hierarchical modeling, achieving significant performance improvements. Similarly, HyperSAT (HyperSAT) accomplishes this through a combination of graph sampling and a key-value joint attention mechanism. We consider hierarchical modeling of KGs to be a promising direction, existing approaches are customized for HKG form and difficult to generalize to other types of facts.
@@
Chung et al. (bive) are the first to introduce nested facts to model relationships between facts, and also propose BiVE which bridges semantics between atomic facts and fact nodes via a simple MLP and scores both atomic facts and nested facts using quaternion-based KGE scoring functions like QuatE (quate) or BiQUE (bique). Based on BiVE, NestE (neste) represents fact nodes using a $1\times 3$ embedding matrix and the nested relations as a $3\times 3$ matrix to avoid information loss. GRADATE (GRADATE) enhances entity and fact representation learning by mining latent intra-fact semantics. However, due to the complexity of NKG representations, existing methods have so far primarily focused on capturing intra-fact semantic information.
@@
Recent studies on TKG representation learning have mainly focused on designing elegant time-aware modules to enhance representation capability. Advanced models like TGeomE+ (TGeomE+) improve the modeling of local semantics in TKGs through 4th-order tensor factorization and linear temporal regularization. Similarly, HGE (HGE) and 5EL (5EL) enhance the expressiveness of temporal spaces by introducing geometric spaces. In contrast, ECEformer (eceformer) captures only inter-fact semantics by unfolding entity neighborhood subgraphs along the temporal chain to model temporal information. However, TKG representation learning methods that can simultaneously capture both intra-fact temporal semantics and global structural information remain largely unexplored. UniHR achieves this by regarding timestamps as nodes and directly employing hierarchical message passing.
@@
Unlike existing triple-based systems (RDF) including RDF (triple) reification, RDF-star and labeled RDF representation techniques, we constrain “triple” to serve as the basic units of HiDR form, and split “nodes” and “relations” into three types respectively, making it more suitable for graph representation learning. Meanwhile, “triple” form makes HiDR could continuously benefit from the model developments of triple-based KGs, which is the most active area of research for link prediction over KGs.
@@
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec (Time2Vec):
@@
NaLP (NaLP) 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
StarE (stare) 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
GRAN (gran) 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
tNaLP (tnalp) 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
HyNT (hynt) 0.479 0.411 0.601 0.478 0.409 0.601 0.328 0.256 0.468 0.356 0.285 0.493
ShrinkE (shrinke) 0.485 0.431 0.601 - - - 0.345 0.275 0.482 - - -
HAHE (hahe) 0.498 0.418 0.610 0.497 0.421 0.614 0.343 0.269 0.484 0.378 0.306 0.515
NYLON (nylon) 0.385 0.299 0.527 0.384 0.300 0.520 0.326 0.262 0.446 0.291 0.226 0.414
HyperSAT (HyperSAT) 0.493 0.427 0.610 0.496 0.430 0.613 0.345 0.270 0.489 0.380 0.306 0.520
@@
Table 1: Results on HKG datasets, are reproduced by us and others are taken from (HyperSAT).
@@
Since the query varies across different settings, we use the transformer (transformer) as the decoder with mask pattern. Specifically, we serialize the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
@@
For HKGs, we use WikiPeople (NaLP) and WD50K (gran). For NKGs, we select FBH, FBHE and DBHE (bive). For TKGs, we use wikidata12k (hyte). To further evaluate the potential of the unified representation, we further introduce hyper-relational TKG datasets WIKI-hy and YAGO-hy (HypeTKG).
@@
We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings (transe) during the evaluation to eliminate existing facts in the dataset. It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set. Similarly, the difference in filtering settings of TKG occurs in timestamp. In this paper, we adopt the strict filtering setting of the former. To ensure fair comparison, for HKG we utilize the results of HyperSAT (HyperSAT) with the same settings as ours. For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
@@
For HKG, we compare with NaLP (NaLP), StarE (stare), GRAN (gran), tNaLP (tnalp), HyNT (hynt), ShrinkE (shrinke), HAHE (hahe), NYLON (nylon) and HyperSAT (HyperSAT). For NKG, QuatE (quate), BiQUE (bique), BiVE (bive), NestE (neste), HOKE (HOKE) and GRADATE (GRADATE) are chosen as baselines. For TKG, we compare against following methods: ATiSE (ATiSE), TGeomE+ (TGeomE+), HGE (HGE), HyperSAT (HyperSAT), DuaTHP (duathp), ECEformer (eceformer) and 5EL (5EL).
@@
QuatE (quate) 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE (bique) 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
BiVE (bive) 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE (neste) 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
HOKE (HOKE) - - - - 3.06 0.719 0.777 2.82 0.674 0.764 2.10 0.674 0.777
GRADATE (GRADATE) - - - - 18.15 0.780 0.871 26.81 0.603 0.757 4.72 0.654 0.916
@@
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from (bive) for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting.
@@
ATiSE (ATiSE) 0.252 0.148 0.288 0.462
TGeomE+ (TGeomE+) 0.333 0.232 0.361 0.546
HGE (HGE) 0.290 0.176 0.323 0.514
ECEformer (eceformer) 0.262 0.159 0.255 0.462
DuaTHP (duathp) 0.304 0.209 0.331 0.509
HyperSAT (HyperSAT) 0.313 0.240 0.348 0.456
5EL (5EL) 0.311 0.237 0.355 0.546
@@
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. Unlike previous methods that use customized decoders for triples, our unified approach does not.
@@
Figure 4: (a) improvements of joint training on hybrid tasks. (b) improvements of joint training on wikimix dataset with hybrid fact forms. Yellow region indicates improvements achieved by joint training.
@@
To analyze the contribution of different modules across various KG types, we present ablation results in Table 4. It can be observed that both intra-fact and inter-fact message passing contribute to performance improvement. In particular, intra-fact message passing proves to be more beneficial for NKGs. We attribute this to fact nodes in NKGs being inherently composed of other atomic nodes, making triple prediction rely heavily on comprehensive bi-level fact semantics. In contrast, HKGs and TKGs focus solely on atomic nodes, whose representations are not dependent on other nodes. Therefore, inter-fact message passing, by capturing the global context among facts, works more effective for HKGs and TKGs, leading to further performance improvements.
@@
Owing to its unified representation, UniHR can flexibly generalize to compositional knowledge graphs, such as hyper-relational temporal KGs (HTKGs) (HypeTKG), which integrate the characteristics of both HKGs and TKGs. In HTKGs, each hyper-relational fact is associated with a timestamp that explicitly indicates its temporal validity. As shown in Table 5, UniHR offers a performance improvement in link prediction tasks on HTKGs, outperforming both TKG-specific and HKG-specific models. This result illustrates the strong ability of UniHR to jointly model auxiliary key-value pairs and temporal information. Furthermore, UniHR achieves competitive performance with the specialized model HypeTKG, despite not relying on complex module stacking.
@@
For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same KG during the message-passing phase under our unified representation form. Therefore, we attempt joint training on two tasks using the NKG dataset, as shown in Fig 4(b). Consistent with previous studies (GRADATE), we also observe that results of joint training are generally superior to those of separate training, further confirming that nested and atomic facts can mutually enhance and complement each other’s semantics.
@@
We believe unified representation is key to develop pre-trained models that integrate multiple KG types. To explore this potential, we jointly train on different KG types. Notably, real-world KGs like Wikidata (wikidata) naturally contain diverse fact types. Thus, we construct a hybrid dataset wikimix, by filtering two Wikidata subsets: HKG WikiPeople and TKG wikidata12k, which share 3,546 entities and 18 relations but have no overlapping facts. To prevent data leakage, we remove test entries whose main triples appear in the other subset’s train set (537 from wikidata12k and 384 from WikiPeople).